OpenStudy (anonymous):

find second derivative of y^(1/2)+xy

OpenStudy (anonymous):

please explain how you arrived at your answer

OpenStudy (amistre64):

i see a power rule, a product rule, and a chain rule ... assuming this is implicits

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im just having a little trouble with a few of my homework problems

OpenStudy (amistre64):

the main issue with implicits is that you are only used to working with y; but forget that x and y mean anything and just consider that operations that relate to them. tell me, how would you think a power rule for y^4 looks like?

OpenStudy (anonymous):

4y^3y'

OpenStudy (amistre64):

very good, and for this one just apply the exponent it uses ... and a product rule for xy ?

OpenStudy (anonymous):

i know that the first derivative should be 1/2y^-(1/2)y'+y+y' right? but I can't figure out the second derivative of this

OpenStudy (amistre64):

\[y^{1/2}+xy\] \[\frac12y^{-1/2}~y'+(x'y+xy')\] solve for y' cause we will need it in the second derivative; and x' wrtx = 1 \[D[\frac12y^{-1/2}~y'+y+xy']\] \[D[\frac12y^{-1/2}~y']+D[y]+D[xy']\] \[\frac12D[y^{-1/2}~y']+D[y]+D[xy']\\power/product~~~~~~~~~~~~~~~product\]

OpenStudy (amistre64):

\[\frac12D[y^{-1/2}~y']+D[y]+D[xy']\] \[\frac12D[y^{-1/2}~y']+y'+x'y'+xy''\] \[\frac12(-\frac12y^{-3/2}~y'y'+y^{-1/2}~y'')+y'+y'+xy''\] \[-\frac14y^{-3/2}~(y')^2+\frac12y^{-1/2}~y''+y'+y'+xy''\] solve for y''

OpenStudy (amistre64):

\[-\frac14y^{-3/2}~(y')^2+\frac12y^{-1/2}~y''+y'+y'+xy''\] \[-\frac14y^{-3/2}~(y')^2+\frac12y^{-1/2}~y''+2y'+xy''\] \[-\frac14y^{-3/2}~(y')^2+y''(\frac12y^{-1/2}+x)+2y'\] \[y''(\frac12y^{-1/2}+x)=\frac14y^{-3/2}~(y')^2-2y'\] \[y''=\frac{\frac14y^{-3/2}~(y')^2-2y'}{\frac12y^{-1/2}+x}\]

OpenStudy (amistre64):

you can simplify it to your hearts content .....

OpenStudy (anonymous):

thanks I'm going to work on this and try to get the hang of it.

OpenStudy (amistre64):

good luck :) i find it simpler to get the form of the second derivative and then fill in the y' parts; but sometimes it might be simpler to simplify the y's first .... personal preferences and all

OpenStudy (anonymous):

i"m just having trouble with all the negative powers and all the rational powers it makes it really confusing especially when I can't just leave them in the form they always want it positive and with the roots. For example the answer the book got was \[\frac{ 2y ^{2}(5+8x \sqrt{y}) }{ (1+2x \sqrt{y})^{3} }\]

OpenStudy (amistre64):

lets determine a suitable expression for y' \[\frac12y^{-1/2}~y'+(x'y+xy')\] \[\frac12y^{-1/2}~y'+y+xy'\] \[y'(\frac12y^{-1/2}+x)=-y\] \[y'=-\frac{y}{\frac12y^{-1/2}+x}\] and sort it \[y'=-\frac{y}{\frac1{2\sqrt{y}}+x}\] \[y'=-\frac{2y\sqrt{y}}{1+2x\sqrt{y}}\] now, we have something to insert into the y' parts

OpenStudy (amistre64):

lets clean up the y'' \[y''=\frac{\frac14y^{-3/2}~(y')^2-2y'}{\frac12y^{-1/2}+x}\] \[y''=\frac{\frac{(y')^2}{4y\sqrt{y}}-2y'}{\frac1{2\sqrt{y}}+x}\] \[y''=\frac{\frac{(y')^2}{2y}-4\sqrt{y}y'}{1+2x\sqrt{y}}\] \[y''=\frac{(y')^2-8y\sqrt{y}y'}{2y(1+2x\sqrt{y})}\] with any luck i mathed that out correctly

OpenStudy (amistre64):

\[y''=\frac{(-\frac{2y\sqrt{y}}{1+2x\sqrt{y}})^2-8y\sqrt{y}(-\frac{2y\sqrt{y}}{1+2x\sqrt{y}})}{2y(1+2x\sqrt{y})}\] \[y''=\frac{\frac{4y^3}{1+2x^2y+4x\sqrt{y}}+\frac{16y^3}{1+2x\sqrt{y}}}{2y(1+2x\sqrt{y})}\] your right, it is a pain :)

OpenStudy (amistre64):

lets assume i did the y' simplification correctly :) \[y'=-\frac{2y\sqrt{y}}{1+2x\sqrt{y}}\] \[y'=-\frac{(1+2x\sqrt{y})(2y\sqrt{y})'-....}{1+2x\sqrt{y}}\]still a pain; lets do a product instead of a quotient

OpenStudy (amistre64):

\[y'=-(2y\sqrt{y})({1+2x\sqrt{y})^{-1}}\] \[y'=-(2yy^{1/2})({1+2xy^{1/2})^{-1}}\] \[y'=-2(y'y^{1/2}+\frac12yy^{-1/2})({1+2xy^{1/2})^{-1}}\] lol .... its just messy no matter what i think of

OpenStudy (amistre64):

\[y''=\frac{\frac{4y^3}{(1+2x\sqrt{y})^2}+\frac{16y^3}{1+2x\sqrt{y}}}{2y(1+2x\sqrt{y})}\] \[y''=2y\frac{\frac{2y^2}{(1+2x\sqrt{y})^2}+\frac{8y^2}{1+2x\sqrt{y}}}{2y(1+2x\sqrt{y})}\] \[y''=\frac{\frac{2y^2}{(1+2x\sqrt{y})^2}+\frac{8y^2}{1+2x\sqrt{y}}}{1+2x\sqrt{y}}\] \[y''=\frac{\frac{2y^2}{(1+2x\sqrt{y})^2}+\frac{8y^2(1+2x\sqrt{y})}{(1+2x\sqrt{y})^2}}{1+2x\sqrt{y}}\] \[y''=\frac{2y^2+8y^2(1+2x\sqrt{y})}{(1+2x\sqrt{y})(1+2x\sqrt{y})^2}\] \[y''=\frac{2y^2(1+4(1+2x\sqrt{y})}{(1+2x\sqrt{y})^3}\] looking closer

OpenStudy (anonymous):

wow thank you so much you my good fellow have gone above and beyond the call of duty

OpenStudy (amistre64):

lol, it was bothering me :) good luck

OpenStudy (amistre64):

if this came up in real life; i would have left it as 2 variable y'' and y'; solving for y' would have created a solid value to apply in the y'' and we could have avoided all the senseless simplifications ....

OpenStudy (anonymous):

most definitely it was one of the last problems on the homework anyway probably not going to be anything near that difficult on a test but i just wanted to know how to do it.